# PromptCodec: High-Fidelity Neural Speech Codec using Disentangled Representation Learning based Adaptive Feature-aware Prompt Encoders

### v1

@@ Abstract
Neural speech codec has recently gained widespread attention in generative speech modeling domains, like voice conversion, text-to-speech synthesis, etc. However, ensuring high-fidelity audio reconstruction of speech codecs under high compression rates remains an open and challenging issue. In this paper, we propose PromptCodec, a novel end-to-end neural speech codec model using disentangled representation learning based feature-aware prompt encoders. By incorporating additional feature representations from prompt encoders, PromptCodec can distribute the speech information requiring processing and enhance its capabilities. Moreover, a simple yet effective adaptive feature weighted fusion approach is introduced to integrate features of different encoders. Meanwhile, we propose a novel disentangled representation learning strategy based on cosine distance to optimize PromptCodec’s encoders to ensure their efficiency, thereby further improving the performance of PromptCodec. Experiments on LibriTTS demonstrate that our proposed PromptCodec consistently outperforms state-of-the-art neural speech codec models under all different bitrate conditions while achieving impressive performance with low bitrates.
@@ 1 Introduction
Specifically, by incorporating a Mel-Spectrogram based conditional prompt encoder and a pre-trained FBank-based voice-print prompt encoder [21], PromptCodec not only enhances the upper limit of its encoder’s capabilities but also diverts the information that needs to be predicted. In this way, PromptCodec can effectively improve the model’s performance at low bitrates. Additionally, to ensure the efficiency of PromptCodec encoders, we propose a novel disentangled representation learning (DRL) based strategy to optimize the proposed PromptCodec. Third, a simple yet effective adaptive feature-weighted fusion (AFWF) method is introduced, which can further enhance the performance of PromptCodec.
@@
  • We present a novel cosine distance based disentangled representation learning based technique to constrain and optimize the encoders of the propose PromptCodec, ensuring their efficiency.
@@ 3 METHODOLOGY
In this section, we first provide an overview of the proposed PromptCodec. Subsequently, the detailed architecture and principles of the proposed prompt encoders, cosine distance based DRL strategy, adaptive feature weighted fusion (AFWF) method, and training loss function are given.
@@ 3.1 Overview
Assume $x\in R^{d\times f}$ denotes the input audio signal, where $d$ and $f$ are the duration and sampling rate of the audio. Our proposed PromptCodec first encodes the input $x$ into the feature representation space $z\in R^{d\times f/M}$ using the conventional codec encoder, where M is the striding factor of the conventional encoder. Then, the GRVQ-based quantizer produces a compressed representation $z_{q}$ via learnable discreted codes. Simultaneously, the Mel-spectrogram $x_{M}$ and FBank $x_{F}$ features of $x$ are fed into the a conditional prompt encoder and a pre-trained voice-print prompt encoder [21], and their corresponding features $z_{PC}$ and $z_{PV}$ are aligned with $z_{q}$ using multi-layer linear projection (MLP) modules. Furthermore, to improve the information utilization efficiency of these encoders, we disentangle the above captured features, i.e., $z_{q}$, $z_{PC}$ and $z_{PV}$, employing a novel cosine distance based disentangled representation learning strategy. Finally, the decoder reconstructs the input audio $\tilde{x}$ from the adaptively weighted features of $z_{q}$, $z_{PC}$ and $z_{PV}$.
@@ 3.2.2 FBank-based voice-print Encoder
3.3 Cosine Distance based DRL Strategy
Employing multiple encoders to distribute the information required for speech reconstruction inevitably leads to information redundancy among them. Hence, to enhance their information utilization efficiency, we propose a novel cosine distance-based DRL learning strategy. Specifically, we calculate the cosine similarity between each pair of these three encoders’ features, and minimize their cosine similarity as an additional penalty added to the final training loss function, forcing them to predict different information. By this means, the encoders of PromptCodec can fulfill their respective roles, thereby improving their information utilization efficiency and enhaning the overall performance of PromptCodec.
@@ 3.5 Training Loss Function
The training loss of PromptCodec comprises five components: reconstruction loss, discriminator loss, feature matching loss, GRVQ commitment loss, and cosine similarity-based DRL loss.
To be specific, for the reconstruction loss $L_{rec}$ and GRVQ commitment loss $L_{vq}$, we follow a similar method to HifiCodec [11]. Regarding the discriminator loss $L_{adv}$ and feature matching loss $L_{f}$, we solely utilize MPD [35], MS-STFTD [10], and L1 feature matching loss [36]. As for the cosine similarity based DRL Loss $L_{DRL}$, it includes the following three parts:
@@
where SSIM denotes the cosine distance based similarity function, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight coefficients to adjust $L_{1}$, $L_{2}$, and $L_{3}$.
@@ 4.2 Main Results
PromptCodec 3.677 0.976 0.568
As illustrated in Table 1, we can easily observe that the proposed PromptCodec significantly outperforms the Encodec, HifiCodec, and TiCodec methods in terms of all metrics under different bitrates. Notably, when at low bitrate, i.e., $n_{q}=1$, PromptCodec achieves 65.1% and 9.7% relative PESQ and STOI improvements and a 34.2% relative MCD reduction over the second best performing TiCodec. From the perspective of PESQ, the proposed PromptCodec demonstrates advantages in speech quality reconstruction, as well as its powerful ability to maintain the fidelity of the original speech, which validates the effectiveness of our proposed prompt encoder enhancement strategy.
@@ 4.3 Ablation Study
4.3.1 Effect of the proposed cosine distance based DRL
First, we evaluate the effectiveness of the proposed cosine distance based DRL strategy. As seen in Table 1, when removing the cosine distance based DRL technique, the performance of PromptCodec consistently drops in terms of all metrics, indicating our design improving the information utilization efficiency of the encoders.

### v2

@@ Abstract
Neural speech codec has recently gained widespread attention in generative speech modeling domains, like voice conversion, text-to-speech synthesis, etc. However, ensuring high-fidelity audio reconstruction of speech codecs under low bitrate remains an open and challenging issue. In this paper, we propose PromptCodec, a novel end-to-end neural speech codec using feature-aware prompt encoders based on disentangled representation learning. By incorporating prompt encoders to capture representations of additional input prompts, PromptCodec can distribute the speech information requiring processing and enhance its representation capabilities. Moreover, a simple yet effective adaptive feature weighted fusion approach is introduced to integrate features of different encoders. Meanwhile, we propose a novel disentangled representation learning strategy based on structure similarity index measure to optimize PromptCodec’s encoders to ensure their efficiency, thereby further improving the performance of PromptCodec. Experiments on LibriTTS demonstrate that our proposed PromptCodec consistently outperforms state-of-the-art neural speech codec models under all different bitrate conditions while achieving superior performance with low bitrates.
@@ 1 Introduction
Specifically, by incorporating a Mel-Spectrogram based conditional prompt encoder and a pre-trained FBank-based voice-print prompt encoder [21], PromptCodec not only enhances the upper limit of its encoder’s capabilities but also diverts the information that needs to be predicted. In this way, PromptCodec can effectively improve the model’s performance at low bitrates. Additionally, to ensure the efficiency of PromptCodec encoders, we propose a novel structure similarity index measure (SSIM) [22] based disentangled representation learning (DRL) strategy to optimize the proposed PromptCodec. Third, a simple yet effective adaptive feature-weighted fusion (AFWF) method is introduced, which can further enhance the performance of PromptCodec.
@@
  • We present a novel SSIM based disentangled representation learning based technique to constrain and optimize the encoders of the propose PromptCodec, ensuring their efficiency.
@@ 3 METHODOLOGY
In this section, we first provide an overview of the proposed PromptCodec. Subsequently, the detailed architecture and principles of the proposed prompt encoders, SSIM based DRL strategy, adaptive feature weighted fusion (AFWF) method, and training loss function are given.
@@ 3.1 Overview
Assume $x\in R^{d\times f}$ denotes the input audio signal, where $d$ and $f$ are the duration and sampling rate of the audio. Our proposed PromptCodec first encodes the input $x$ into the feature representation space $z\in R^{d\times f/M}$ using the conventional codec encoder, where M is the striding factor of the conventional encoder. Then, the GRVQ-based quantizer produces a compressed representation $z_{q}$ via learnable discrete codes. Simultaneously, the Mel-spectrogram $x_{M}$ and FBank $x_{F}$ features of $x$ are fed into the a conditional prompt encoder and a pre-trained voice-print prompt encoder [21], and their corresponding features $z_{PC}$ and $z_{PV}$ are aligned with $z_{q}$ using multi-layer linear projection (MLP) modules. Furthermore, to improve the information utilization efficiency of these encoders, we disentangle the above captured features, i.e., $z_{q}$, $z_{PC}$ and $z_{PV}$, employing a novel SSIM based disentangled representation learning strategy. Finally, the decoder reconstructs the input audio $\tilde{x}$ from the adaptively weighted features of $z_{q}$, $z_{PC}$ and $z_{PV}$.
@@ 3.2.2 FBank-based voice-print Encoder
3.3 SSIM based DRL Strategy
Employing multiple encoders to distribute the information required for speech reconstruction inevitably leads to information redundancy among them. Hence, to enhance their information utilization efficiency, we propose a novel SSIM based DRL learning strategy. Specifically, we calculate the SSIM between each pair of these three encoders’ features, and minimize their SSIM as an additional penalty added to the final training loss function, forcing them to predict different information. By this means, the encoders of PromptCodec can fulfill their respective roles, thereby improving their information utilization efficiency and enhaning the overall performance of PromptCodec.
@@ 3.5 Training Loss Function
The training loss of PromptCodec comprises five components: reconstruction loss, discriminator loss, feature matching loss, GRVQ commitment loss, and SSIM based DRL loss.
To be specific, for the reconstruction loss $L_{rec}$ and GRVQ commitment loss $L_{vq}$, we follow a similar method to HifiCodec [11]. Regarding the discriminator loss $L_{adv}$ and feature matching loss $L_{f}$, we solely utilize MPD [36], MS-STFTD [10], and L1 feature matching loss [37]. As for the SSIM based DRL Loss $L_{DRL}$, it includes the following three parts:
@@
where SSIM denotes the SSIM based similarity function, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight coefficients to adjust $L_{1}$, $L_{2}$, and $L_{3}$.
@@ 4.2 Main Results
PromptCodec 3.720 0.976 0.561
As illustrated in Table 1, we can easily observe that the proposed PromptCodec significantly outperforms the Encodec, HifiCodec, and TiCodec methods in terms of all metrics under different bitrates. Notably, when at low bitrate, i.e., $n_{q}=1$, PromptCodec achieves 64.0% and 9.7% relative PESQ and STOI improvements and a 35.6% relative MCD reduction over the second best performing TiCodec. From the perspective of PESQ, the proposed PromptCodec demonstrates advantages in speech quality reconstruction, as well as its powerful ability to maintain the fidelity of the original speech, which validates the effectiveness of our proposed prompt encoder enhancement strategy.
@@ 4.3 Ablation Study
4.3.1 Effect of the proposed SSIM based DRL
First, we evaluate the effectiveness of the proposed SSIM based DRL strategy. As seen in Table 1, when removing the SSIM based DRL technique, the performance of PromptCodec consistently drops in terms of all metrics, indicating our design improving the information utilization efficiency of the encoders.